# Repairs in a Block World: A New Benchmark for Handling User Corrections with Multi-Modal Language Models

### v1

@@ 9 Potential Risks
Appendix A BlockWorld-Repairs
@@
A.1 AMT Collection
Each task or Human Intelligence Task (HIT) in AMT consisted of 4 dialogues, with an initial easier, training dialogue that workers were free to experiment with. AMT workers could repeat the task for as long as there were block configurations they had not seen before. The instructions are in Figure 4, with the user interface of the task itself in Figure 5. We introduced many attention checks, particularly with the repairs, where workers were asked to select whether the robot had performed the correct action (Yes/No) before entering free-text repairs. If their answer did not match the expected answer too many times, they would get a warning and be limited from submitting any subsequent HITs.
@@ A.1 AMT Collection
A.2 Human Baselines
Participants interacted with the interface shown in Figure 6, where the chat was already filled by a randomly selected dialogue. Participants had to read the instructions and repairs and provide their best interpretation of which block the instructions referred to and where to place it. Once satisfied with their choice, they would confirm and proceed to the next dialogue until the session ended. This interface uses the Unity engine (v2022.3) [^6] https://unity.com/ to generate images based on the original Block World configurations.
@@ A.2 Human Baselines
A.3 Dataset Statistics
Table 5 provides additional statistics about the dataset.
@@ A.3 Dataset Statistics
Table 5: Summary of Collected Entries
@@
Appendix B Experimental Setup
B.1 Prompts
@@
Target Position Prediction Answer only with the final bounding box location mentioned. The bounding box consists of 4 values between 0 and 1. Here is an example: [0.000, 0.123, 0.075, 0.204]
@@
B.2 Model Fine-Tuning
@@
Appendix C Additional Error Analysis
@@
Table 9: Mean and Standard Deviation (SD) for test dialogues in BW-R by difficulty level and task.
Appendix D Additional Experimental Results

### v2

@@ 9 Potential Risks
Acknowledgements
Chiyah-Garcia’s PhD is funded under the EPSRC iCase with Siemens (EP/T517471/1). We thank the anonymous reviewers for their insightful comments that helped improve this paper. Finally, the authors acknowledge the use of the Heriot-Watt University high-performance computing facility (DMOG) and associated support services in the completion of this work.
Appendix A Related Datasets
We compare BlockWorld-Repairs with related works in Table 5. Our work builds on a robotic manipulation task involving common pick&place actions and rich referring expressions in a situated dialogue. Similar to our benchmark, both TEACh Padmakumar et al. (2022) and DialFRED Gao et al. (2022) extend the popular ALFRED Shridhar et al. (2020) to include dialogues with clarificational exchanges. From a high-level ALFRED directive (e.g., “prepare coffee”), these works crowd-source possible questions for low-level actions that may be underspecified in the original task (e.g., “where is the coffee mug?”). Thus, they often operate on the last level of Clark’s (1996) joint action ladder, action and consideration, where the goal is already clear and lack of knowledge (underspecified location of objects or actions available) represents the primary ambiguity (e.g., VIMA-Bench; Jiang et al., 2023)). These works rarely feature bi-directional information flow ($human\leftrightarrows agent$) in their dialogues, unlike BW-R which ensures dialogue dependency with partial robot movements and candidate responses. DialFRED, Alexa Arena Gao et al. (2023) and CoDraw iCRs Madureira and Schlangen (2023) contain questions for underspecified actions/objects (e.g., “Which spoon should I pick up?”) but, in these situations, the ambiguity is reduced to a minimum so that candidates are easily distinguishable with strong visual representations (e.g., “yellow or blue spoon?” or “left or right?”).
By contrast, SIMMC 2.0 Kottur et al. (2021) includes a large number of visually ambiguous objects (e.g., 5 identical red t-shirts in view) within long dialogues that reference multiple items (averaging 4.5±2.4 unique objects per dialogue). This makes it rich in cross-modal coordination phenomena, with terse referring expressions that mix spatial, historical and visual cues (e.g., “Pick the red shirt. Which one? The one on the right wardrobe, above the blue jumper”). Similarly, Cups Dobnik et al. (2020) aligns with our work, as it also explores coordination in highly ambiguous environments where many objects share the same shape/colour but are in different positions, requiring complex referring expressions. However, Cups focuses on Frame of Reference coordination rather than manipulation or resolving coreferences. BW-R is aimed at exploring this coordination as part of a collaborative human-robot tabletop manipulation task.
Related Work Manipulation Ref Exp Dialogue Repairs Miscommunication
[…]
Appendix B BlockWorld-Repairs
@@
B.1 AMT Collection
Each task or Human Intelligence Task (HIT) in AMT consisted of 4 dialogues, with an initial easier, training dialogue that workers were free to experiment with. AMT workers could repeat the task for as long as there were block configurations they had not seen before. The instructions are in Figure 4, with the user interface of the task itself in Figure 5. We introduced many attention checks, particularly with the repairs, where workers were asked to select whether the robot had performed the correct action (Yes/No) before entering free-text repairs. If their answer did not match the expected answer too many times, they would get a warning and be limited from submitting any subsequent HITs. We used the CRWIZ framework Chiyah Garcia et al. (2020) to control many of these details, including the robot’s non-verbal actions (i.e., movements, pointing), effectively running a simulation or sequence of events different to each worker in AMT.
@@ B.1 AMT Collection
B.2 Human Baselines
Participants interacted with the interface shown in Figure 6, where the chat was already filled by a randomly selected dialogue. Participants had to read the instructions and repairs and provide their best interpretation of which block the instructions referred to and where to place it. Once satisfied with their choice, they would confirm and proceed to the next dialogue until the session ended. This interface uses the Unity engine (v2022.3) [^7] https://unity.com/ and De Pellegrin and Petrick (2024) to generate environments based on the original Block World configurations.
@@ B.2 Human Baselines
B.3 Dataset Statistics
Table 6 provides additional statistics for the BlockWorld-Repairs dataset.
@@ B.3 Dataset Statistics
Table 6: Summary of collected entries in BW-R.
@@
Appendix C Experimental Setup
C.1 Prompts
@@
Target
Position
Prediction
Answer only with the final bounding box location mentioned. The bounding box consists of 4 values between 0 and 1. Here is an example: [0.000, 0.123, 0.075, 0.204]
@@
C.2 Model Fine-Tuning
@@
Appendix D Additional Error Analysis
@@
We also provide in Table 11 the results of selecting GPT-4o performance as the proxy for example difficulty. Since we only have 1 prediction value for each BW-R entry and we cannot do the average as with the human annotations, we instead order the data by distance (source and target separately) and divide it into three equal sizes at the 33rd and 66th percentiles: 1) Easy category contains the top 33% of entries (shortest distance to source/target); 2) Medium contains entries that are between the 33% and 66% distances; and 3) Hard is reserved for the worst performing 33% entries.
We find that Idefics2 and LLaVA have low performance on all levels, whilst GPT-4o only surpasses humans on easy dialogues (not surprising), but quickly deteriorates to distances worse than random on hard dialogues. Interestingly, humans perform well across all difficulties except the hard target predictions. We also observe that the mean word trends get stronger (Table 10), although we leave the analysis to prove a correlation for future work. These results further show that models do not process TPRs as consistently as humans.
@@
GPT-4o Performance
Easy 29.95 (±12.09) 34.29 (±13.81)
Medium 29.62 (±12.02) 31.46 (±12.45)
Hard 26.52 (±12.31) 24.54 (±11.73)
Table 10: Mean and Standard Deviation (SD) for test dialogues in BW-R by difficulty level and task. We differentiate between using human performance (Section 6) and GPT-4o performance (Appendix D) as the difficulty proxies.
Difficulty Model Source Target
Acc ↑ Mean (SD) ↓ Mean (SD) ↓
Easy Idefics2 0.43 2.60 (±2.6) 6.11 (±2.9)
LLaVA 0.48 2.02 (±2.0) 4.20 (±2.9)
Humans 0.78 1.07 (±1.1) 2.05 (±2.1)
GPT-4o 0.93 .00 (±.0) 1.27 (±.5)
Medium Idefics2 0.47 2.29 (±2.3) 4.97 (±2.6)
LLaVA 0.47 2.01 (±2.0) 4.25 (±2.7)
Humans 0.77 1.05 (±1.1) 2.23 (±2.3)
GPT-4o 0.58 1.26 (±1.3) 2.63 (±.4)
Hard Idefics2 0.04 4.91 (±4.9) 5.92 (±3.1)
LLaVA 0.22 4.61 (±4.6) 5.69 (±3.4)
Humans 0.64 2.52 (±2.5) 4.84 (±3.7)
GPT-4o 0.00 8.33 (±8.3) 6.31 (±3.1)
Table 11: Model performances across subsets with increasing levels of complexity (easy, medium, hard) using GPT-4o performance as the difficulty proxy.
Appendix E Additional Experimental Results
@@
Appendix F Other Notes
Object detection and predictions
Splitting the task into object detection and then prediction out of a list of candidates did not show any advantages in our initial tests, possibly because these models are already pretrained with object detection objectives Liu et al. (2023); Laurençon et al. (2024). We already see better performance with source block prediction than with target location prediction. Thus, we unified both tasks into a single common task format (bounding box prediction) for easier comparison. Additionally, we would like to point out that target prediction would be a harder task because it requires the model to derive a bounding box for an empty board location, which models particularly struggle as shown in Section 6.
Repairs facilitate the task
Repairs are fundamental capabilities in human conversations, which our human study further shows with the increasing performance with TPRs (see Table 3). TPRs have the dual role of narrowing down the candidate pool (ie., “not this block”) and providing new information to uniquely identify the correct referent (ie., “just left of there”). Our paper shows that current out-of-the-box models do not have the capabilities to exploit repairs to facilitate the task (unlike humans), but we should try to distil these skills with alternative training regimes that facilitate learning from interaction data. Handling TPRs is essential for real-world human-robot collaboration in novel scenarios and this paper explores this gap between models and humans.